OpenStudy (anonymous):

For what intervals is g(x) = 1 ---------- x^2 + 1 concave down? _____ < x < _________ please explain? thanks!!

OpenStudy (anonymous):

so it would be -∞ < x < ∞ ? :/

OpenStudy (anonymous):

@eliassaab your reply disappeared! :/

OpenStudy (anonymous):

\[g(x)=\frac{1}{x^2+1}\\ g'(x)=-\frac{2x}{(x^2+1)^2}\\ \begin{align*}g''(x)&=\frac{-2(x^2+1)^2-(-2x)(2(x^2+1))(2x)}{(x^2+1)^4}\\&=\frac{-2(x^2+1)^2+8x^2(x^2+1)}{(x^2+1)^4}\\ &=\frac{(x^2+1)\left(-2(x^2+1)+8x^2\right)}{(x^2+1)^4}\\ &=\frac{6x^2-2}{(x^2+1)^3}\end{align*}\] Find the critical points (when is \(g''(x)=0\) or undefined?).

OpenStudy (anonymous):

Here is the answer it concave down for \[ -\frac 1{\sqrt 3}\le x \le\frac 1{\sqrt 3} \]

OpenStudy (anonymous):

This where the second derivative is negative

OpenStudy (anonymous):

so you simplified 6x^2-2/(x^2+1)^3 into + or - 1/√3 ? :/ if so, could you please show me how to simplify that? @SithsAndGiggles i'm following what you've written so far:)

OpenStudy (anonymous):

The sigh depends on the numerator sign \[ 6x^2 -2\]

OpenStudy (anonymous):

SIGN

OpenStudy (anonymous):

Now \[ 6 x^2 -2 = 6( x - \frac1{\sqrt 3})( x + \frac1{\sqrt 3}) <0\\for \\ -\frac 1{\sqrt 3}\le x \le\frac 1{\sqrt 3} \]

OpenStudy (anonymous):

The denominator of the second derivative is always positive. So we need to study the sign of the numerator only.

OpenStudy (anonymous):

Is it clear now?

OpenStudy (anonymous):

yes :) thanks so much!! :D

OpenStudy (anonymous):

YW